OpenStudy (marigirl):

Could you please explain mathematically what actually happens to the absolute value below

OpenStudy (marigirl):

\[\ln \left| 2k^3-1 \right|=7.5\]

OpenStudy (marigirl):

Then I would go \[\left| 2k^3-1 \right|=e^(7.5)\]

OpenStudy (marigirl):

what will happen to the absolute function value now?

OpenStudy (irishboy123):

try\[2k^3−1=\pm e^{7.5}\]

OpenStudy (marigirl):

My question was to solve for k which the value is + or - 9.67 how do we 'get rid' of an absolute value function, or how do we balance an absolute value function in an equation Since absolute value function only look at positive values, I am a little confused with your response @IrishBoy123

OpenStudy (irishboy123):

let's say |x| = 3, then we can conclude that x = 3 OR x = -3 yes?

OpenStudy (irishboy123):

ie, in that example i give, we conclude \[x = \pm3\]

OpenStudy (irishboy123):

so if \(|x|=e^{7.5}\), you must consider \(x= \pm e^{7.5}\)

OpenStudy (irishboy123):

you should get two answers \[\sqrt[3]{\frac{1 \pm e^{7.5}}{2}}\]

OpenStudy (marigirl):

oh yea it makes sense now, I got k=9.67 and k=-9.66 Thanks heaps! I really appreciate it

OpenStudy (irishboy123):

\(\checkmark\)